A parallel plate capacitor has an area of 5cm² and 0.1mm apart. IN between the plates is a sheet of mica ( 4.8x10¯¹¹C²/N.m² ). Find its capacitance.


1
Expert's answer
2022-03-02T10:07:01-0500
"C=\frac{\epsilon A}{d}\\
=\frac{4.8*10^{-11}*5*10^{-4}}{0.1*10^{-3}}=2.4*10^{-10}\:\rm F"